### v2

# Generating coupled cluster code for modern distributed memory tensor software

Abstract
Using GPU-based HPC platforms efficiently for coupled cluster computations is a challenge due to heterogeneous hardware structures. The constant need to adapt software to these structures and the required man-hours makes a systematization of high-performance code development desirable, even more so for higher-order coupled cluster. This is generally achieved by introducing a high-level representation of the problem, which is then translated to low-level instructions for the hardware using a compiler/translator component. Designing such software comes with another challenge: Allowing efficient implementation by capturing key symmetries of tensors, while retaining the abstraction from the hardware. We review ways to address these two challenges while presenting design decisions which led us to the development of a general-order coupled cluster code generator. The systematically produced code shows excellent weak scaling behavior running on up to 1200 GPUs using the distributed memory tensor library ExaTENSOR. We present an open-source modular tensor framework "tenpi" for coupled cluster code development with diagrammatic derivation, visualization module, symbolic algebra, intermediate optimization and support for multiple tensor backends. Tenpi brings higher-order CC functionality to the massively parallel ExaCorr module of the DIRAC code for relativistic molecular calculations.
I INTRODUCTION
According to Jack Dongarra, the cofounder of the TOP500 list,TOP500.org (2024) 99% of the FLOP performance of modern supercomputers lies in GPU accelerators.Dongarra (2023); Chirigati (2022) However, in the domain of distributed coupled cluster calculations, the focus of the present work, most implementations are able to use only about 10% of the theoretical maximum FLOP rate.Calvin et al. (2021)
This exposes the difficult situation in which scientists find themselves.Ma, Krishnamoorthy, and Agrawal (2011); Woolston (2022) Implementing fixed BLAS (Basic Linear Algebra Subroutines) and MPI (Message-Passing Interface) statements is no longer sufficient as modern machines have heterogeneous structures.Lyakh (2019) Fig. 1 contains a schematic representation of a node of the Summit supercomputer. Communication throughputs and memory sizes differ between the components by orders of magnitude. Restructuring data movement across multiple levels of hierarchy to reduce the cost is a challenging problem,Ma, Krishnamoorthy, and Agrawal (2011) where the developer basically faces a graph theoretical task.Thibault (2018) The traditional 5-year lifespan of supercomputersRojas et al. (2019) has accelerated to about 3 years due to AI-driven breakthroughs in energy efficiencyZhao et al. (2023) and keeping up with the changes requires expert manpower for which academia competes with industry.Woolston (2022); Heffernan and McKay (2018)
The situation benefits systematic approaches to development and parallelization.Ma, Krishnamoorthy, and Agrawal (2011); Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) One way is to build the code such that it can be quickly adapted to underlying numerical software changes using a code generator.Lechner et al. (2024); Papadopoulos (2019); Auer et al. (2006) Another way is to rely on tools which allow a high level of abstraction, like MATLAB,The Mathworks, Inc. (2024) MapleWaterloo Maple, Inc. (2024) or Mathematica,Wolfram Research, Inc. (2024) to write in parallel frameworks like SYCLThe Khronos Group, Inc. (2021) and OpenACC,The OpenACC Organization (2022) to use runtime environments like StarPU,Augonnet et al. (2011) PaRSEC,Herault et al. (2021) MADNESS,Harrison et al. (2016) and LEGION,Bauer et al. (2012) or tensor compilers like DISTAL.Yadav, Aiken, and Kjolstad (2022)
In any case, one tries to i) abstract from hardware-specific code by using a high-level representation of the problem: creation/annihilation operator strings,Lechner et al. (2024) their expansion using Wick’s theorem,Bochevarov, Valeev, and David Sherrill (2004) diagrams,Shiozaki et al. (2008) tensor operations,Lotrich et al. (2009) portable tasksAugonnet et al. (2011) and data distribution schemes.Yadav, Aiken, and Kjolstad (2022) ii) The software then translates this to instructions for hardware.
The challenge is to allow efficient implementation by capturing key mathematical features and symmetries of the problem, while retaining the abstraction from the hardware.Matthews (2018a) Strong candidates for finding this balance are tensor compilersYadav, Aiken, and Kjolstad (2022) which have seen an explosion of interestZheng et al. (2024) in the computer science (CS) community. However, to enable production use in the coupled cluster domain, performance-critical features like index permutation symmetry or block sparsity would need to be implemented.Panyala (2014); Shiozaki et al. (2008) The same holds for major machine learning libraries. Unlike in the case of AI, where big tech companies drive the development, in coupled cluster and tensor networks, scientists are left to rely on themselves. This was concluded at a panel discussion with representatives of these companies at a recent CECAM workshopBrandejs et al. (2024) which we co-organized.
The purpose of this paper is to explore the path from a high-level representation of CC methods to an efficient code on modern machines from the viewpoint of software-architecture. We provide the reader with a state-of-the art on the necessary tensor software, and we benchmark tenpiBrandejs and Saue (2024) - our solution to the two challenges: i) adapting CC to heterogeneous HPC, ii) identifying a balanced high-level representation.
I.1 Elegant development of CC methods, systematic approach
Equations in coupled cluster theory can reach a degree of complexity where manual manipulation is no longer practical.Kállay and Surján (2001) Higher orders of the method with quadruple excitations and above include hundreds of terms with many indices. This accuracy is required to reach the scale of quantum-electrodynamics (QED) effects which our group is invested in.Saue et al. (2024)
Even though derivation by hand can be simplified by using diagrams which help detect equivalent terms, subsequent work with many indices is error-prone when done by hand. This and further manipulations like the design of intermediates with respect to a given cost-functionLai et al. (2012) or implementing other flavors of the method are clear candidates for automatic symbolic treatment.
There have been numerous efforts in this direction with differing degrees of success. Here we focus on complete toolchains which include both equation and code generation - these are often linked to an existing quantum chemistry package. A prominent example and one of the pioneering systematization projects is the Tensor Contraction Engine (TCE)Baumgartner et al. (2005) initially developed by Hirata,Auer et al. (2006) a part of NWCHEM software.Kendall et al. (2000) Even though TCE was not the first such code generator,Janssen and Schaefer (1991); Li and Paldus (1994); Kállay and Surján (2001); Nooijen and Lotrich (2001) to the best of our knowledge, it has the most complete description in literature.Lechner et al. (2024) Later works often follow its scheme:Shiozaki et al. (2008)
1.
Derivation (Derive the formulas of the method.)
2.
Optimization (Optimize the expressions to reduce computational complexity.)
3.
Transformation (Map to binary tensor contractions of a math library.)
Another noteworthy abstraction effort is the domain-specific language (DSL) SIAL (superinstruction assembly language) for tensor operations in the ACES III package.Lotrich et al. (2009); Deumens et al. (2011) Even though initially successful, its limitation was the same as that of TCE: customizability. Not all methods from the CC family can be efficiently expressed in the form of tensor contractions and without C++ or Fortran frontend support, one could not easily include custom code and optimizations.Panyala (2014) Some key parallelization optimizations even had to be done by hand in final Fortran loops of the generated code, as it was too difficult to modify the generator itself.Ozog et al. (2013) Further packages using their own tensor DSL include QChemShao et al. (2006) based on the libtensorEpifanovsky et al. (2013) library, and Cyclops CTFSolomonik et al. (2013) whose performance for distributed CPU contractions gained distinction in the computer science community, where it is used as a reference.Calvin, Lewis, and Valeev (2015); Yadav, Aiken, and Kjolstad (2022); Irmler et al. (2023)
CC formula generators themselves divide in three main groups based on their approach to the derivation: i) by applying the algebra of creation and annihilation operators, ii) by Wick’s theorem, iii) or by diagrams.
Unlike TCE, which was based on Wick’s theorem, the derivation inside the MRCC package of Kállay and SurjánKállay and Surján (2001) uses a representation of Kucharski-Bartlett diagramsShavitt and Bartlett (2009); Kucharski and Bartlett (1992) based on strings of integers. In terms of derivation, the present paper builds on their approach.
The SMITH generator of Shiozaki Shiozaki et al. (2008); Shiozaki (2017) is based on antisymmetrized Goldstone diagrams.Goldstone (1957); Kucharski and Bartlett (1992) SMITH comes with an elegant input format in terms of second-quantized operators sandwiched between Slater determinants. Compared to its predecessors, SMITH is able to produce intermediates with index permutation symmetry for a broader class of methods.Shiozaki et al. (2008) The latest version 3 has transitioned from Wick’s theorem to second-quantized operators and was used to generate parts of the BAGEL package.Park and Shiozaki (2017)
Recent ORCANeese (2022) is good example of a large-scale systematization effort where a substantial part of the package is generated using the ORCA-AGE generatorLechner et al. (2024) based directly on application of the algebra of creation and annihilation operators. Even though this approach is very general, it took up to 7 years before performance issues with the generation were resolved.Krupička et al. (2017); Lechner et al. (2024)
New efforts have emerged in past years, as groups behind quantum chemical packages consider having their own toolchain. Most of them use Wick’s theorem: i) SeQuant of the Valeev group Bochevarov, Valeev, and David Sherrill (2004) which is a C++ rewrite of their Mathematica code, ii) Drudge/Gristmill of the Scuseria groupSong, Henderson, and Scuseria (2022) which can switch between different abstract algebras and where the use of Wick’s theorem is optional, iii) FEMTO of SaitowSaitow, Kurashige, and Yanai (2013) used for DMRG-MRCI and pair-natural orbitals, iv) SQANeuscamman, Yanai, and Chan (2009), v) ebccBackhouse (2024) and vi) GeCCo.Köhn, Richings, and Tew (2008) These efforts can be further classified by supported methods and features, see more detailed reviews.Lechner et al. (2024); Solomonik et al. (2014) Many other efforts exist which focus on only one of the three steps of the process.G. A. Smith and Gray (2018); Quintero-Monsebaiz and Loos (2023); Lyakh, Ivanov, and Adamowicz (2004); Li and Paldus (1994); Wang, Peyton, and Crawford (2022)
I.2 Optimization step
Algorithmically the most complicated part of the entire process is the automatic design of intermediates. Finding the optimal set of intermediates is NP-hard,Chi-Chung, Sadayappan, and Wenger (1997); Lechner et al. (2024) so heuristics is used in practice. The ideal cost function would be the actual walltime of the target calculation, which depends on the size of the application system, the hardware and the underlying math library. As this is unknown a priori, most optimizers rely on a naive FLOP count cost model, despite the fact that distributed CC calculations are often communication-bound, meaning that this model does not take into account the leading term of the walltime. Most chemistry codes are limited only to the basic single-term contraction path optimization and do not implement global (multi-term) optimization, which is replaced by hand-tuning.
The TCE also relies on search-based approach driven by a FLOP count performance model, using a global optimizer OpMin written by Sadayappan et al.Lai et al. (2012) The optimization consists of:
1.
Single term optimization (contraction path within a term)
2.
Factorization (distributive law)
3.
Common subexpression elimination (reuse equivalent terms)
This is a common structure among tensor contraction optimizers, with differing level of sophistication.Panyala (2014) OpMin performs well compared with hand-tuned code of NWCHEM for higher-order CC, also because its cost-function takes into account the index-permutation symmetry. Nevertheless, it does not support perturbative approaches with energy denominators.
Other efforts include a distributed GPU-contraction optimizer benchmarked for CCSD(T),Ma, Krishnamoorthy, and Agrawal (2011) followed by the AutoHOOT optimizerMa, Ye, and Solomonik (2020) (not CC), and an optimizer with a GPU-aware cost model.Harju et al. (2013) The aforementioned complete toolchains also include own optimizers of varying levels of sophistication. For more information, see the literature overview in the thesis of PanyalaPanyala (2014) on loop-level optimization.Gao et al. (2006); Sahoo et al. (2005) Note that there is no clear boundary on how low-level an optimizer should be. If it reaches low-level, it can be called a compiler, which is in fact a trend in state-of-the-art works.Yadav, Aiken, and Kjolstad (2022); Yadav et al. (2024); Zheng et al. (2024) There is a strong connection between tensor compilers and compiler optimizations for different applications.Johnson et al. (2001); Meirom et al. (2022); Bondhugula et al. (2008) However, CC contractions require specific treatment due to their distinctive features:Panyala (2014)
1.
A specific index permutation symmetry
2.
Fully permutable for loops (order of summations)
3.
Dependencies preventing loop fusion never occur
Last but not least, a noteworthy effort is the load balancer DLTCLai et al. (2013) which analyzes dependencies between contractions and groups them in layers that can be executed concurrently. This is particularly relevant for higher-order CC with large number of contractions with widely different computational cost.Solomonik et al. (2014)
I.3 State-of-the-art: Tensor contraction
Tensor contraction represents the most computer-intensive operation of numerous methods in quantum chemistry, condensed matter physics, nuclear physics, machine learning and quantum computing.Psarras et al. (2022); Matthews (2018b); Brabec et al. (2020) Examples include coupled cluster methods,Calvin et al. (2021); Peng et al. (2016); Herault et al. (2021) tensor networks,Levy, Solomonik, and Clark (2020) quantum computing simulators,Lyakh et al. (2022) certain neural networks Sharir, Shashua, and Carleo (2022) and signal processing methods.Chi and Kolda (2012); Sidiropoulos et al. (2017) Common to all these is that the key limitation to the affordable system size is the cost of tensor contraction.
Typically, the software packages decompose tensor contractions into primitive matrix operations and pass them to BLAS.Shi et al. (2016); Georganas et al. (2021) There are different approaches of dealing with tensor transposition (reshape), which is in general required for the usage of BLAS GEMM (General Matrix Multiplication), which depends on a unit-stride index to multiply two matrices. Given a tensor memory layout, an index is said to have a unit stride when incrementing it by one translates to a move by one scalar in memory. Note that in the context of tensor operations, the CS community uses the term transposition interchangeably with reshape. The most common approach is TTGTMatthews (2018b) (Transpose-Transpose-GEMM-Transpose): for
1.
Transpose and into unit-stride form
2.
Use GEMM to execute the contraction
3.
Transpose result array to obtain
In his block-scatter matrix tensor contraction (BSMTC) scheme, MatthewsMatthews (2018b) has shown that is possible to hide the transposition cost by loading tensor parts into CPU L2 and L3 cache in order which effectively performs the required permutation of its elements.
Another alternative is the GETT scheme (GEMM-like tensor-tensor multiplication) by Springer and BientinesiSpringer and Bientinesi (2017) which generates code to call matrix-matrix multiplication kernels while again reformulating the sub-matrix-packing. A GPU implementation thereof has become a foundation of the single-node library cuTENSOR.cuT (2019)
Regarding the parallelization strategy, most software relies on OpenMP+MPI parallelization on homogenous CPU-based computational clusters.Calvin et al. (2021) The advent of GPU-based supercomputers has rendered this paradigm obsolete.Herault et al. (2021)
Comparing to the software for matrix operations, we currently identify two major practical hurdles for tensor software:
1.
Contrary to BLAS for matrix operations, there is no standardized interface for tensor operations. This causes substantial code duplication.Psarras et al. (2022)
2.
Contrary to CPU-based software,Solomonik et al. (2013); Peng et al. (2016) there is no established GPU-based implementation of a tensor contraction library with support for features required by the community such as distributed memory and block sparsity,Herault et al. (2021) which would offer sufficient level of maintenance and optimization for current supercomputers.
As for standardization, there has been some work in the past, but none of the interfaces have prevailed. BTASShi et al. (2016) has aimed at providing a standard and a basic CPU implementation. TBLISMatthews (2018b) provides BLAS-like tensor calls. Until today, there is no GPU implementation of these. During the 2022 Dagstuhl tensor workshop,Bientinesi et al. (2022) a development of domain-specific tensor language has been initiated, but remains unfinished and undocumented. There has been work-technical specifications of additional aspects of standardization, like tensor-memory distributionValeev (2023); cppreference.com community (2023) and randomized multilinear algebra.Murray et al. (2023) Currently, there are recurring meetings across academia and private sector under the TAPP (Tensor Algebra Processing Primitives) standardization initiative.TAPP working group (2024)
Regarding the implementations, there are numerous scattered efforts listed in Ref. Psarras et al., 2022. However, for open-source distributed memory GPU libraries, there are only a few major players. i) ExaTENSORLyakh (2019); Pototschnig et al. (2021) based on TAL-SHtal (2018), where the key developer has left to industry. ii) TiledArrayPeng et al. (2016) and Cyclops CTF,Solomonik et al. (2013) where the solid support for GPUs is still under development. iii) TACO DISTAL,Yadav, Aiken, and Kjolstad (2022) which has performance issues for higher-order tensors. iv) TAMMMutlu et al. (2023) scales well on supercomputers, but has a potentially problematic dependency of Global Arrays,Palmer et al. (2022) with a small user base for a communication library. v) cuTENSORcuT (2019); Zhang et al. (2020) is proprietary and only supports Nvidia hardware, while vi) hipTensorMillette, Ma, and Karunanidhi (2024) is at an early stage of development. Other industry products like vii) PyTorchPaszke et al. (2019) and viii) TensorFlowAbadi (2016) focus almost exclusively on machine learning and again lack features required in CC.
II THEORY AND IMPLEMENTATION
II.1 Coupled cluster amplitude equations and their generation
The coupled cluster method is based on an exponential ansatz
where is an excitation operator, and the corresponding cluster amplitude tensor. in our case denotes the Hartree–Fock (HF) reference.
Even though the code generator is designed to go to arbitrary order and in practice generates optimized code up to CCSDTQP (see Table 1), in the presented applications we restrict ourselves to up to quadruple excitations
We use and to denote occupied and virtual orbitals respectively.
After defining the similarity transformed normal-ordered Hamiltonian as , we can write down the CC energy equations
which determine the energy and the cluster amplitudes.
The explicit equation terms can then be derived using different techniques. In our implementation, we chose the diagrammatic scheme of Kállay and SurjánKállay and Surján (2001) (see the reasoning in Section II.2). This general-order coupled cluster derivation scheme is based on a string representation of CC diagrams.
As depicted in Fig. 2, an integer string of length 13 is used to represent a diagram. Two diagrams are equivalent up to a sign if their strings are the same. There is a simple algorithm depicted in Fig. 3 that uses this representation to generate distinct diagrams for each excitation level of the CC Eqs. (3).
To find the explicit equation terms, tenpiBrandejs and Saue (2024) translates the integer strings into line graph representation, i.e. an edge list of particle and hole lines connecting operators and external bra indices of Eqs. (3). Such a representation makes the application of coupled cluster interpretation rules similar to when done by hand. All used high-level representations are summarized in Fig. 5 and explained in Sections II.2 and II.3.
II.2 Design decisions
The initial goal of the tenpi project was to create a programming environment for relativistic coupled cluster which separates science from the computational platform by getting tensor developments under the hood. This has three aspects:
1.
Systematic development of higher-order coupled cluster methods which include hundreds of tensor contractions.
2.
Development without having to consider the parallelization strategy.
3.
Automatic treatment of tensor symmetries, such as index-permutation symmetry and block-sparsity (for systems with spatial symmetry).
This paper addresses the first two points and prepares the ground for the third point.
The design choices are limited by practical considerations. The toolchain should not be restricted to a given flavor of the coupled cluster method, but should be written in a general way and be able to go to high orders of the theory. It should allow manipulations of the method by PhD and Master students of quantum chemistry. This led to the decision to base the formula generator on Kucharski-Bartlett diagrams,Shavitt and Bartlett (2009); Kucharski and Bartlett (1992) since diagrams are visual and thus more accessible for students than the other CC derivation schemes listed in section I.1. The ease of use for target users is one of the reasons why we chose python to implement the generator, aside from the fact that it is suitable for text processing and symbolic operations. The string representation of CC diagrams of Kállay and SurjánKállay and Surján (2001) was chosen because it clearly distinguishes equivalent diagrams and can be used to generate distinct connected diagrams (see section II.1) in an intuitive way.Kállay and Surján (2001)
The diagram strings are next translated into directed graphs, by storing a list of edges (adjacency matrix). This representation makes the implementation of the CC interpretation rules readable. Advanced users can define a custom set of rules (customizability) or for instance extend the Diagram class to support multiple interaction operators and other generalizations, without having to change the rest of the code (modularity).
To follow closely the hand-typeset style of the pedagogical texts of Refs. Crawford and Schaefer, 2000 and Shavitt and Bartlett, 2009, tenpi prints the diagrams in PDF format using the CCDiag LaTeXpackage.Kats (2024) During the development of higher-order CC methods this visual representation proved itself invaluable for debugging. Overall, the emphasis on intuitive, readable, modular and visual implementation makes problems directly visible and modules unit-testable (see section II.3).
The workflow of the tenpi code generator is depicted in Fig. 4 and corresponding high-level representations in Fig. 5. A key step is to optimize the expressions obtained. Since writing a global optimizer from scratch is beyond the scope of this work, we interfaced our code with OpMin, which was chosen over GristmillSong, Henderson, and Scuseria (2022) due to a favorable performance for higher-order CC.Lai et al. (2012) However, even though the heuristics of OpMin is reported to stay within of the global minimum FLOP-count, the number of intermediates it generates tends to be too costly in terms of memory. Higher-order CC tensors in our calculations cannot be stored on disk due to prohibitive I/O costs. Our approach is to distribute them among the CPU and GPU RAM memory of the nodes. This comes with a limitation in the amount of available distributed memory due to communication overhead growing quadratically with the number of nodes in the round-robin distribution scheme used (see below). Therefore, we implemented our own secondary optimization, which rolls back part of the intermediates by plugging them back in their respective terms. The number of intermediates is reduced by about a factor of three at the expense of increasing the FLOP count (see Fig. 6). This approach is justified as long as the calculation stays communication-dominated. We do not reach a perfect balance, which would require setting up a communication-based cost-function. Optimization of this aspect is the subject of ongoing work.
During the optimization, we originally represented contractions using the SymPy packageMeurer et al. (2017) but later we turned to the use of a custom representation of tensor contractions for greater flexibility (Contraction class, see section II.3). Even though SymPy is widely used, its current user interface is cumbersome for higher-order tensors.
The accessibility of the code should not come at the cost of performance. Therefore, the generated production code is strictly in a compiled language, in our case Fortran 2008 in order to integrate smoothly with the ExaCorr modulePototschnig et al. (2021) of the DIRAC package.Visscher et al. (2024); Saue et al. (2020)
As we aim for scalability on modern GPU exascale machines, we turned to the use of a tensor library to execute the final tensor kernels (ExaTENSOR, TAL-SH). ExaTENSOR was chosen from the limited list of available distributed GPU tensor libraries (see section I.3) due to its excellent performance for relativistic CCSD.Pototschnig et al. (2021) In the present work, the performance is reproduced and demonstrated to scale to up to 1200 GPUs by automatically generated code (see Fig. 9).
II.3 Implementation
An equation term is implemented using the ‘Contraction’ class, which includes a list of tensors, a scalar factor and optionally permutation operators and an output tensor. Contraction class instances are collected in a list, which forms the left-hand side of Eqs. (3).
Tensor class includes a list of upper and lower indices. Each index is an Index class instance. Index class has its type (e.g. occupied, virtual, active virtual), a convention according to which it is printed (e.g. ; ) and then a number corresponding to the order in which the specific character comes in the class (e.g. 1, 2, 1). Types and conventions are easily customizable by modifying a single location in the code, which greatly facilitates the implementation of methods and symmetries with multiple index spaces or tiling.Mutlu et al. (2023) Index, Tensor and Contraction class are fully compatible with smart functionality to account for block-sparsity in tensors and to keep track of index permutation symmetry in intermediates explicitly. Work is in progress to include this functionality in the production code. A set of equations can switch between conventions with a single command. This is important because tenpi supports multiple input and output formats to interface with ExaTENSOR, TAL-SH, OpMin, NumPy, including a possibility to input equations by hand (see Fig. 5). For the sake of modularity, all these codes are integrated by extending one of two predefined interfaces: GetCodeInterface or ParseCodeInterface. These specify how to implement a new format of output or input respectively. Thanks to this design it is straightforward to interface tenpi with a new program.
The optimization of intermediates is an error-prone procedure and checking the results by hand can be extremely tedious. Therefore, following the example of OpMin, tenpi includes verification of correctness using a generated python brute-force test code with contractions expressed in nested for loops. Random tensors with tiny virtual and occupied sizes are used to check whether the optimized equation produces the same result as the original one.
| method | CCD | CCSD | CCSDT | CCSDTQ | CCSDTQP |
| number of diagrams | 11 | 48 | 102 | 183 | 289 |
| derivation step | 0.03 s | 0.1 s | 0.3 s | 0.6 s | 0.8 s |
| OpMin | 0.03 s | 40 s | 22 min | 7 h 10 min | 4 days 1 h |
| 2nd optimization | 0.02 s | 0.1 s | 16 s | 1 h 9 min | 17 days |
| code generation | 0.01 s | 0.1 s | 3 s | 2 min | 2 h 13 min |
The derivation step of tenpi is very fast, see timings in Tab. 1. We have verified by comparing with Ref. Kállay and Surján, 2001 that tenpi generates the correct number of diagrams for each of the CC Eqs. (3) up to excitation level 20. Even though it generates a highly-parallel code, tenpi itself is written as sequential python. Despite this, the most demanding step of global intermediate optimization still performs quite well.
Aside from amplitude equations, tenpi can already generate matrix elements (see Figs. 3, 5, cf. SMITHShiozaki et al. (2008); Shiozaki (2017)) and is currently being extended to generate response density matrices.
When going from CCD and CCSD levels during the development, an error in energy appeared. It was found quickly thanks to the visual representation of diagrams: For the CC diagram interpretation rules to yield a correct sign, e.g. for the case of amplitude-equation projected onto , the line starting in has to end in , and not in . This rule is often assumed implicitly.Shavitt and Bartlett (2009); Crawford and Schaefer (2000) Transition from CCSDT and CCSDTQ proved to be more peculiar, due to a mistake in an algorithm for unrolling index permutations. The problem was unraveled once a brute force unit test was written for the corresponding module.
III COMPUTATIONAL DETAILS
III.1 Machines used
The calculations were performed on the machines listed in Table 2 which includes their technical specification. When ExaTENSOR is applied for coupled cluster,Pototschnig et al. (2021) the most relevant parameters for the performance are the interconnect bandwidth, RAM (random-access) and HBM (high-bandwidth) memory, as well as GPU specifications. Over the two supercomputer generations represented in Table 2 the bandwidth and the HBM have grown about fourfold, while RAM has basically stagnated. Despite the rapid development of GPU processing power, the table shows that the evolution of other important parameters has been relatively slow over the last 5 years. Recently, large memory CPU nodes have emerged which go against this trend, e.g. Intel Xeon 6900P with MRDIMM technology (Multiplexed Rank Dual In-line Memory Module) and 3 TB of fast memory as announced for an HPC platform in Japan.NEC (2024)
We would like to bring to the attention of the reader that one should not expect an easy installation and tuning of distributed memory tensor libraries, as it mostly requires GPU and MPI experts. ExaTENSOR features cutting-edge parallelization strategy with multithreading, nested OpenMP threads, GPU-aware MPI and one-sided communication, which proved itself difficult to support for some HPC platforms. For instance on Frontier, an internal resource starvation within Cray MPICHHewlett Packard Enterprise (2023) (HPE) causes certain MPI requests to be blocked, and the calculation hangs as a result. The engineers from the supplier company (HPE) are currently working on the issue.
| machine | Frontier | LUMI | Karolina | Summit | Olympe |
| laboratory | OLCF | CSC data center | IT4Innovations | OLCF | CALMIP |
| organization | U.S. DOE | EuroHPC | Cz. edu. ministry | U.S. DOE | CNRS |
| GPU nodes | 9408 | 2978 | 72 | 4600 | 48 |
| GPUs per node | 8 | 8 | 8 | 6 | 4 |
| GPU type | AMD MI250x | AMD MI250x | NVIDIA A100 | NVIDIA V100 | NVIDIA V100 |
| HBM per GPU | 64 GB | 64 GB | 40 GB | 16 GB | 16 GB |
| CPU type | AMD EPYC 7713 | AMD EPYC 7A53 | AMD EPYC 7452 | IBM POWER9 | Intel Skylake |
| cores per node | 64 | 64 | 64 | 168 | 36 |
| RAM per node | 512 GB | 512 GB | 1024 GB | 512 GB | 384 GB |
| network | Cray Slingshot | Cray Slingshot-11 | Infiniband HDR200 | Infiniband EDR | Infiniband EDR |
| net. bandwidth | 4x 25 GB/s | 4x 25 GB/s | 4x 6.25 GB/s | 23 GB/s | 12.5 GB/s |
III.2 Relativistic Hamiltonian
The state-of-the-art post-HF four-component relativistic calculations employ the no-pair approximation.Sucher (1980); Almoukhalalati et al. (2016) This approximation is well-substantiated for applications of chemical scale. Formally, it yields a Hamiltonian analogous to the non-relativistic case
where the indices only run over the positive-energy spinors that span the one-electron basis.
In the present work we employ, unless otherwise stated, the exact two-component (X2C) Hamiltonian,Jensen (2005); Kutzelnigg and Liu (2005); Iliaš and Saue (2007); Liu and Peng (2009) where the summations in Eq. (4) are limited to positive energy orbitals by construction. Relativistic two-electron picture-change corrections were added either using the AMFI packageHess et al. (1996); Schimmelpfennig or using the recently implemented amfX2C (XAMFI) correction.Knecht et al. (2022) The Gaunt/Breit terms are not included in the present work.
III.3 Geometry, basis, etc.
For the benchmarks, we used the UF6 molecule in dyall.v2z basisDyall (2023) with the X2C Hamiltonian, and bond distance of Å. In the corresponding ground-state energy CC calculations, we used a HF reference. Despite the high Oh symmetry, there is no related performance gain in the CC calculations as the implementation does not support symmetry yet.
Furthermore, we studied CO in cc-pVTZ basisDunning (1989) at bond-distances between and Å. There, we used the amfX2C correction for two-electron scalar-relativistic and spin-orbit picture-change effects arising within the X2C Hamiltonian framework.Knecht et al. (2022)
For the cases where the full virtual space is not used, we then calculated the MP2 frozen natural orbitalsYuan, Visscher, and Gomes (2022) (with a cutoff on occupation number) on top of which the final CC calculation was done.
We performed two correctness checks with a nonrelativistic Hamiltonian. For the first test, we chose H2O in a 3-21G basisBinkley, Pople, and Hehre (1980) at the equilibrium internuclear distance of Å and H-O-H angle of . The second test is with LiH in 6-31G basisDill and Pople (1975); Ditchfield, Hehre, and Pople (1971) at the internuclear distance of Å. Both CC correctness tests were performed with respect to the HF reference.
All the CC calculations were performed in uncontracted basis sets. When referring to sizes of occupied and virtual spaces, we use the notation of AS(), where is the number of electrons and is the number of Kramers pairs active in a coupled cluster calculation (occupied+virtual). One Kramers pair corresponds to two spinors related by time reversal symmetry. In all calculations, the choice of occupied and virtual spaces respects the rising orbital energies or the occupation numbers in the case of MP2. See the used software in Section IV.1.
IV RESULTS AND DISCUSSION
IV.1 Correctness: H2O and LiH
To demonstrate the correctness, we compare ground state coupled cluster energies from tenpi (present work) with the established MRCC packageKállay et al. (2020); mrc (2024); Kállay and Surján (2001); Kállay, Gauss, and Szalay (2003) and with existing handwritten code in DIRACVisscher et al. (2024); Saue et al. (2020) where possible, i.e. either with the initial DIRAC RelCC moduleVisscher, Lee, and Dyall (1996) or with the more recent massively parallel ExaCorr module.Pototschnig et al. (2021) The correctness check was performed on two small systems: H2O (Table 3) and LiH (Table 4) and shows an agreement within the convergence error.
| method | code | convergence555the final error shown in the run | Total energy [] |
| SCF | DIRAC SCF | 1.1E-12111energy difference | -75.58 5498 7542 |
| SCF | MRCC | 4.4E-13111energy difference | -75.58 5498 7680 |
| MP2 | DIRAC ExaCorr | -75.66 6586 2130 | |
| MP2 | MRCC | -75.66 6586 2218 | |
| CCSD | DIRAC tenpi444using the TAL-SH library | 0.5E-09333norm of the residual vector | -75.67 2947 6512 |
| CCSD | MRCC | 9.6E-10333norm of the residual vector | -75.67 2947 6563 |
| CCSDT | DIRAC tenpi222using the ExaTENSOR library | 0.4E-09333norm of the residual vector | -75.67 3875 1038 |
| CCSDT | MRCC | 4.9E-10333norm of the residual vector | -75.67 3875 1087 |
| CCSDTQ | DIRAC tenpi222using the ExaTENSOR library | 0.1E-08333norm of the residual vector | -75.67 3974 4768 |
| CCSDTQ | MRCC | 3.5E-10333norm of the residual vector | -75.67 3974 4817 |
| method | code | convergence444the final error shown in the run | Total energy [] |
| SCF | DIRAC SCF | 1.4E-12111energy difference | -7.97 9321 5650 |
| SCF | MRCC | 6.2E-15111energy difference | -7.97 9321 5634 |
| MP2 | DIRAC ExaCorr | -7.99 1935 0613 | |
| MP2 | MRCC | -7.99 1935 0593 | |
| CCSD | DIRAC tenpi222using the ExaTENSOR library | 0.6E-09333norm of the residual vector | -7.99 8346 9438 |
| CCSD | MRCC | 2.9E-10333norm of the residual vector | -7.99 8346 9410 |
| CCSDT | DIRAC tenpi222using the ExaTENSOR library | 0.5E-09333norm of the residual vector | -7.99 8358 3476 |
| CCSDT | MRCC | 4.2E-10333norm of the residual vector | -7.99 8358 3449 |
| CCSDTQ | DIRAC tenpi222using the ExaTENSOR library | 0.3E-05333norm of the residual vector | -7.99 8358 3478 |
| CCSDTQ | MRCC | 7.0E-10333norm of the residual vector | -7.99 8358 3630 |
IV.2 Benchmark: UF6 CCSD
To benchmark the parallel scaling, we compare the performance of CCSD generated by tenpi with hand-tuned code in the ExaCorr module of DIRAC and with generated code from the codegen (mb-autogen) generator.Gomes and Lyakh (2023) First we perform a strong-scaling benchmark on UF6 with fixed problem size. In all cases, the correct total energy of is retrieved with residual norm less than . The parallel speedup for strong scaling benchmark is calculated in a standard way as a walltime ratio for processing units (in this case we consider entire nodes). If is not available due to large memory requirements, the timing per node with the minimal number of nodes is used to estimate it.
The benchmark was performed on the Summit supercomputer of the Oak Ridge Leadership Computing Facility (OLCF) using the GPU implementation of the ExaTENSOR library. Each Summit node is equipped with 6 GPUs (see Table 2). We used up to 300 GPUs simultaneously for strong scaling.
As shown in Fig. 7, the code scales well until about 20 nodes for UF6. There, the curve leaves its approximately linear behavior. This is expected for fixed problem size. As the number of workers increase, they start to suffer from work starvation, which eventually causes the overhead from the unnecessarily increased internode communication to dominate. We chose to model this using the Amdahl’s lawAmdahl (2013) for strong scaling
where is the number of nodes, and are the proportions of serial and parallel code. From fitting, we found , corresponding to when the processes wait for communication during tensor operations and for synchronization barriers which are enforced after each tensor operation.
As expected, the automatically generated code is slightly less efficient than the hand-tuned code. The timings in Fig. 8 show a fixed-factor slowdown of about 30% for tenpi, whereas about 80% for codegen. The former is a good result when compared with ORCA, where 100% timing overhead was a practical rule of thumb threshold for acceptance of generated code. Lechner et al. (2024)
To investigate the influence of the internode communication, we then moved to analyze the weak scaling of the code. Using the same system and method as previously, we varied the size of virtual space such that the number of operations per node is kept fixed, based on expected CCSD scaling of with the number of occupied and of virtual spinors. The weak scaling efficiency is again calculated as . The ideal weak scaling efficiency would be a constant 1.
As demonstrated in Fig. 9, the code exhibits a relatively constant weak scaling behavior to up to 200 nodes, meaning that it can be used efficiently on up to 1200 GPUs. This is a pleasing result for a code without hand-tuning. We attribute most fluctuations in the graph to differing appropriateness of fixed block size of elements used in case of CCSD when distributing the tensors, where is the number of indices. Some virtual sizes come with smaller number of padding zeroes than others. Moreover, the handwritten code has different intermediates than tenpi and the advantage can differ for different space sizes. One could profile and hand-tune the generated code based on several test runs, but such non-systematic intervention would go against the design principles. Instead, using measured walltimes to improve the cost model in the optimization step is a subject of further study.
IV.3 Application: CO
The CO molecule was chosen due to the interesting aspects of its triple bond stretching, studied in recent works of the Shabaev groupUsov et al. (2024) and Koput.Koput (2024) In the former work, the influence of quantum electrodynamics (QED) corrections is analyzed which is relevant for the HAMP-vQED projectSaue et al. (2024) for which the present code was created. These are minute effects and their study requires highly accurate methods.
Fig. 10 shows the ground state energies up to the CCSDT level for the bond distance stretching from to Å. Fig. 11 shows similar results for up to CCSDTQ level with a smaller space.
As expected, single-reference methods struggle at large bond distances. The nonparallelity of CCSD and CCSDT becomes apparent at the bond distance of Å (Fig. 10) and the CC has convergence issues from Å on. This is also visible for the rightmost two points for CCSDTQ in Fig. 11.
A deeper study of the behavior of higher-order CC upon dissociation, and especially analyzing the minimum CC excitation level to break the bond successfully is beyond the scope of this manuscript, and will be a subject for our further work, once index-permutation symmetry of tensors is fully implemented.
V CONCLUSIONS
tenpi is an open-source code generator published under BSD-3 license.Brandejs and Saue (2024) tenpi addresses the complex challenge of implementing coupled cluster for modern GPU based HPC platforms in a software-architecture sound way.
The code generated from tenpi is the first working implementation of CCSDT and CCSDTQ in DIRAC. Excellent weak scaling behavior is demonstrated with up to 1200 GPUs. For massively parallel CCSD, the performance is comparable to the hand-tuned code with the measured overhead of only 30%, which is an improvement over the 80% overhead of the existing generated CCSD. The code generation and optimization itself is fast relative to similar toolchains up to quadruple level.
Overall, the tenpi framework provides users with an elegant and systematic way of implementing the coupled cluster methods while hiding under the hood most of the complexity related to the equations, parallelization and intermediates. A simple python interface is featured, while the output is a highly optimized Fortran code that controls the parallel tensor library.
The stretching of the triple bond of CO was studied with convergence problems observed for higher-order CC, particularly for CCSDTQ at larger bond distances.
The outlook for tenpi includes the development of CCSD(T)Raghavachari et al. (2013) and CCSDT(Q)Kowalski and Piecuch (2000) for distributed-GPU platforms, and addressing the current limitations like the lack of spatial-symmetry support and integration of DIRAC with tensor software that supports index-permutation symmetry. Inclusion of internode communication and real-world performance data into the cost model for intermediate optimization will be studied. The method development can be boosted by a symbolic interface to Mathematica. tenpi can also serve as a standalone production CC package, if the integral generation is interfaced directly to an existing code, like ReSpect.Repisky et al. (2020)
Besides the above, this work maps the latest developments in relevant tensor software and shows how is the transition to a systematic development pipeline necessary to sustain modern coupled cluster codes.
Acknowledgements.
We very much acknowledge the consultations and support from André S.P. Gomes, Dmitry Lyakh, Jiří Pittner, Marcus Wagner, Lucas Visscher, Gabriele Fabbro and Paolo Bientinesi. This project has received funding from the the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101019907). This research used resources of the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725. This work was granted access to the HPC resources of CALMIP supercomputing center under the allocation 2023-p13154 and 2024-M24070. We acknowledge VSB – Technical University of Ostrava, IT4Innovations National Supercomputing Center, Czech Republic, for awarding this project access to the LUMI supercomputer, owned by the EuroHPC Joint Undertaking, hosted by CSC (Finland) and the LUMI consortium through the Ministry of Education, Youth and Sports of the Czech Republic through the e-INFRA CZ (grant ID: 90254). We acknowledge the use of the MRCC packageKállay et al. (2020); mrc (2024) and its CC methods.Kállay and Surján (2001); Kállay, Gauss, and Szalay (2003)References
- TOP500.org (2024) TOP500.org, “TOP500 list,” https://www.top500.org (2024), [Online; accessed 06-Sep-2024].
- Dongarra (2023) J. Dongarra, “A Not So Simple Matter of Software,” https://www.hpcwire.com/2022/11/16/jack-dongarra-a-not-so-simple-matter-of-software/ (2023).
- Chirigati (2022) F. Chirigati, Nature Computational Science 2, 211–212 (2022).
- Calvin et al. (2021) J. A. Calvin, C. Peng, V. Rishi, A. Kumar, and E. F. Valeev, Chemical Reviews 121, 1203 (2021).
- Ma, Krishnamoorthy, and Agrawal (2011) W. Ma, S. Krishnamoorthy, and G. Agrawal, “Practical Loop Transformations for Tensor Contraction Expressions on Multi-level Memory Hierarchies,” in Compiler Construction (Springer Berlin Heidelberg, 2011) p. 266–285.
- Woolston (2022) C. Woolston, “Why science needs more research software engineers,” http://dx.doi.org/10.1038/d41586-022-01516-2 (May 31 2022).
- Lyakh (2019) D. I. Lyakh, International Journal of Quantum Chemistry 119, 25926 (2019).
- Thibault (2018) S. Thibault, On Runtime Systems for Task-based Programming on Heterogeneous Platforms, Habilitation à diriger des recherches, Université de Bordeaux (2018).
- Rojas et al. (2019) E. Rojas, E. Meneses, T. Jones, and D. Maxwell, in 2019 31st International Symposium on Computer Architecture and High Performance Computing (SBAC-PAD) (IEEE, 2019).
- Zhao et al. (2023) D. Zhao, S. Samsi, J. McDonald, B. Li, D. Bestor, M. Jones, D. Tiwari, and V. Gadepally, in Proceedings of the 2023 ACM Symposium on Cloud Computing, SoCC ’23 (ACM, 2023).
- Heffernan and McKay (2018) T. A. Heffernan and A. McKay, Professional Development in Education 45, 102–113 (2018).
- Sum (2024) “Summit User Guide,” https://docs.olcf.ornl.gov/systems/summit_user_guide.html (2024), [Online; accessed 27-May-2024].
- Lechner et al. (2024) M. H. Lechner, A. Papadopoulos, K. Sivalingam, A. A. Auer, A. Koslowski, U. Becker, F. Wennmohs, and F. Neese, Physical Chemistry Chemical Physics 26, 15205–15220 (2024).
- Papadopoulos (2019) A. Papadopoulos, Necessity of Code Generators in Electronic-Structure Theories and their Infrastructure, Master’s thesis, University of Amsterdam, Amsterdam, Netherlands (2019).
- Auer et al. (2006) A. A. Auer, G. Baumgartner, D. E. Bernholdt, A. Bibireata, V. Choppella, D. Cociorva, X. Gao, R. Harrison, S. Krishnamoorthy, S. Krishnan, C.-C. Lam, Q. Lu, M. Nooijen, R. Pitzer, J. Ramanujam, P. Sadayappan, and A. Sibiryakov, Molecular Physics 104, 211–228 (2006).
- The Mathworks, Inc. (2024) The Mathworks, Inc., “MATLAB,” https://mathworks.com/products/matlab.html (2024).
- Waterloo Maple, Inc. (2024) Waterloo Maple, Inc., “Maple,” https://www.maplesoft.com/products/Maple/ (2024).
- Wolfram Research, Inc. (2024) Wolfram Research, Inc., “Mathematica, Version 14.0,” https://www.wolfram.com/mathematica (2024), Champaign, IL.
- The Khronos Group, Inc. (2021) The Khronos Group, Inc., “SYCL,” https://www.khronos.org/sycl/ (2021).
- The OpenACC Organization (2022) The OpenACC Organization, “OpenACC,” https://www.openacc.org (2022).
- Augonnet et al. (2011) C. Augonnet, S. Thibault, R. Namyst, and P.-A. Wacrenier, Concurrency and Computation: Practice and Experience Euro-Par 2009 best papers, 23, 187 (2011).
- Herault et al. (2021) T. Herault, Y. Robert, G. Bosilca, R. J. Harrison, C. A. Lewis, E. F. Valeev, and J. J. Dongarra, in 2021 IEEE International Parallel and Distributed Processing Symposium (IPDPS) (IEEE, 2021).
- Harrison et al. (2016) R. J. Harrison, G. Beylkin, F. A. Bischoff, J. A. Calvin, G. I. Fann, J. Fosso-Tande, D. Galindo, J. R. Hammond, R. Hartman-Baker, J. C. Hill, J. Jia, J. S. Kottmann, M.-J. Yvonne Ou, J. Pei, L. E. Ratcliff, M. G. Reuter, A. C. Richie-Halford, N. A. Romero, H. Sekino, W. A. Shelton, B. E. Sundahl, W. S. Thornton, E. F. Valeev, A. Vazquez-Mayagoitia, N. Vence, T. Yanai, and Y. Yokoi, SIAM Journal on Scientific Computing 38, S123–S142 (2016).
- Bauer et al. (2012) M. Bauer, S. Treichler, E. Slaughter, and A. Aiken, in 2012 International Conference for High Performance Computing, Networking, Storage and Analysis (IEEE, 2012).
- Yadav, Aiken, and Kjolstad (2022) R. Yadav, A. Aiken, and F. Kjolstad, in Proceedings of the 43rd ACM SIGPLAN International Conference on Programming Language Design and Implementation (ACM, 2022).
- Bochevarov, Valeev, and David Sherrill (2004) A. D. Bochevarov, E. F. Valeev, and C. David Sherrill, Molecular Physics 102, 111–123 (2004).
- Shiozaki et al. (2008) T. Shiozaki, M. Kamiya, S. Hirata, and E. F. Valeev, Physical Chemistry Chemical Physics 10, 3358 (2008).
- Lotrich et al. (2009) V. Lotrich, N. Flocke, M. Ponton, B. A. Sanders, E. Deumens, R. J. Bartlett, and A. Perera, in Proceedings of the 23rd international conference on Supercomputing, ICS ’09 (ACM, 2009).
- Matthews (2018a) D. A. Matthews, Molecular Physics 117, 1325–1333 (2018a).
- Zheng et al. (2024) L. Zheng et al., “Awesome Tensor Compilers,” https://github.com/merrymercy/awesome-tensor-compilers (2024).
- Panyala (2014) A. Panyala, Search-based Model-driven Loop Optimizations for Tensor Contractions, Ph.D. thesis, Louisiana State University Libraries (2014).
- Brandejs et al. (2024) J. Brandejs, P. Bientinesi, L. Visscher, A. Gomes, and T. Saue, “CECAM Workshop on Tensor Contraction Library Standardization,” https://tensor.sciencesconf.org/ (2024).
- Brandejs and Saue (2024) J. Brandejs and T. Saue, “tenpi: Tensor Programming Interface,” https://gitlab.com/janbrandejs/tenpi (2024).
- Kállay and Surján (2001) M. Kállay and P. R. Surján, The Journal of Chemical Physics 115, 2945–2954 (2001).
- Saue et al. (2024) T. Saue et al., “HAMP-vQED,” https://dirac.ups-tlse.fr/hamp-vqed/doku.php?id=public:overview (2024).
- Lai et al. (2012) P.-W. Lai, H. Zhang, S. Rajbhandari, E. Valeev, K. Kowalski, and P. Sadayappan, Procedia Computer Science 9, 412–421 (2012).
- Baumgartner et al. (2005) G. Baumgartner, A. Auer, D. Bernholdt, A. Bibireata, V. Choppella, D. Cociorva, X. Gao, R. Harrison, S. Hirata, S. Krishnamoorthy, S. Krishnan, C.-C. Lam, Q. Lu, M. Nooijen, R. Pitzer, J. Ramanujam, P. Sadayappan, and A. Sibiryakov, Proceedings of the IEEE 93, 276–292 (2005).
- Kendall et al. (2000) R. A. Kendall, E. Aprà, D. E. Bernholdt, E. J. Bylaska, M. Dupuis, G. I. Fann, R. J. Harrison, J. Ju, J. A. Nichols, J. Nieplocha, T. Straatsma, T. L. Windus, and A. T. Wong, Computer Physics Communications 128, 260–283 (2000).
- Janssen and Schaefer (1991) C. L. Janssen and H. F. Schaefer, Theoretica Chimica Acta 79, 1–42 (1991).
- Li and Paldus (1994) X. Li and J. Paldus, The Journal of Chemical Physics 101, 8812–8826 (1994).
- Nooijen and Lotrich (2001) M. Nooijen and V. Lotrich, Journal of Molecular Structure: THEOCHEM 547, 253–267 (2001).
- Deumens et al. (2011) E. Deumens, V. F. Lotrich, A. S. Perera, R. J. Bartlett, N. Jindal, and B. A. Sanders, “The Super Instruction Architecture,” in Annual Reports in Computational Chemistry (Elsevier, 2011) p. 179–191.
- Ozog et al. (2013) D. Ozog, J. R. Hammond, J. Dinan, P. Balaji, S. Shende, and A. Malony, in 2013 42nd International Conference on Parallel Processing (IEEE, 2013) p. 30–39.
- Shao et al. (2006) Y. Shao, L. F. Molnar, Y. Jung, J. Kussmann, C. Ochsenfeld, S. T. Brown, A. T. Gilbert, L. V. Slipchenko, S. V. Levchenko, D. P. O’Neill, R. A. DiStasio Jr, R. C. Lochan, T. Wang, G. J. Beran, N. A. Besley, J. M. Herbert, C. Yeh Lin, T. Van Voorhis, S. Hung Chien, A. Sodt, R. P. Steele, V. A. Rassolov, P. E. Maslen, P. P. Korambath, R. D. Adamson, B. Austin, J. Baker, E. F. C. Byrd, H. Dachsel, R. J. Doerksen, A. Dreuw, B. D. Dunietz, A. D. Dutoi, T. R. Furlani, S. R. Gwaltney, A. Heyden, S. Hirata, C.-P. Hsu, G. Kedziora, R. Z. Khalliulin, P. Klunzinger, A. M. Lee, M. S. Lee, W. Liang, I. Lotan, N. Nair, B. Peters, E. I. Proynov, P. A. Pieniazek, Y. Min Rhee, J. Ritchie, E. Rosta, C. D. Sherrill, A. C. Simmonett, J. E. Subotnik, H. Lee Woodcock III, W. Zhang, A. T. Bell, A. K. Chakraborty, D. M. Chipman, F. J. Keil, A. Warshel, W. J. Hehre, H. F. Schaefer III, J. Kong, A. I. Krylov, P. M. W. Gill, and M. Head-Gordon, Phys. Chem. Chem. Phys. 8, 3172–3191 (2006).
- Epifanovsky et al. (2013) E. Epifanovsky, M. Wormit, T. Kuś, A. Landau, D. Zuev, K. Khistyaev, P. Manohar, I. Kaliman, A. Dreuw, and A. I. Krylov, Journal of Computational Chemistry 34, 2293–2309 (2013).
- Solomonik et al. (2013) E. Solomonik, D. Matthews, J. Hammond, and J. Demmel, in 2013 IEEE 27th International Symposium on Parallel and Distributed Processing (IEEE, 2013).
- Calvin, Lewis, and Valeev (2015) J. A. Calvin, C. A. Lewis, and E. F. Valeev, in Proceedings of the 5th Workshop on Irregular Applications: Architectures and Algorithms, SC15 (ACM, 2015).
- Irmler et al. (2023) A. Irmler, R. Kanakagiri, S. T. Ohlmann, E. Solomonik, and A. Grüneis, “Optimizing Distributed Tensor Contractions Using Node-Aware Processor Grids,” in Euro-Par 2023: Parallel Processing (Springer Nature Switzerland, 2023) p. 710–724.
- Shavitt and Bartlett (2009) I. Shavitt and R. J. Bartlett, Many-Body Methods in Chemistry and Physics: MBPT and Coupled-Cluster Theory (Cambridge University Press, 2009).
- Kucharski and Bartlett (1992) S. A. Kucharski and R. J. Bartlett, The Journal of Chemical Physics 97, 4282–4288 (1992).
- Shiozaki (2017) T. Shiozaki, WIREs Computational Molecular Science 8, 1331 (2017).
- Goldstone (1957) J. Goldstone, Proceedings of the Royal Society of London. Series A. Mathematical and Physical Sciences 239, 267–279 (1957).
- Park and Shiozaki (2017) J. W. Park and T. Shiozaki, Journal of Chemical Theory and Computation 13, 3676–3683 (2017).
- Neese (2022) F. Neese, WIREs Computational Molecular Science 12 (2022), 10.1002/wcms.1606.
- Krupička et al. (2017) M. Krupička, K. Sivalingam, L. Huntington, A. A. Auer, and F. Neese, Journal of Computational Chemistry 38, 1853–1868 (2017).
- Song, Henderson, and Scuseria (2022) R. Song, T. M. Henderson, and G. E. Scuseria, The Journal of Chemical Physics 156, 104105 (2022).
- Saitow, Kurashige, and Yanai (2013) M. Saitow, Y. Kurashige, and T. Yanai, The Journal of Chemical Physics 139, 044118 (2013).
- Neuscamman, Yanai, and Chan (2009) E. Neuscamman, T. Yanai, and G. K.-L. Chan, The Journal of Chemical Physics 130, 124102 (2009).
- Backhouse (2024) O. Backhouse, “ebcc,” https://github.com/BoothGroup/ebcc (2024).
- Köhn, Richings, and Tew (2008) A. Köhn, G. W. Richings, and D. P. Tew, The Journal of Chemical Physics 129, 201103 (2008).
- Solomonik et al. (2014) E. Solomonik, D. Matthews, J. R. Hammond, J. F. Stanton, and J. Demmel, Journal of Parallel and Distributed Computing 74, 3176–3190 (2014).
- G. A. Smith and Gray (2018) D. G. A. Smith and J. Gray, Journal of Open Source Software 3, 753 (2018).
- Quintero-Monsebaiz and Loos (2023) R. Quintero-Monsebaiz and P.-F. Loos, AIP Advances 13, 085035 (2023).
- Lyakh, Ivanov, and Adamowicz (2004) D. I. Lyakh, V. V. Ivanov, and L. Adamowicz, The Journal of Chemical Physics 122, 024108 (2004).
- Wang, Peyton, and Crawford (2022) Z. Wang, B. G. Peyton, and T. D. Crawford, Journal of Chemical Theory and Computation 18, 5479–5491 (2022).
- Chi-Chung, Sadayappan, and Wenger (1997) L. Chi-Chung, P. Sadayappan, and R. Wenger, Parallel Processing Letters 07, 157–168 (1997).
- Ma, Ye, and Solomonik (2020) L. Ma, J. Ye, and E. Solomonik, in Proceedings of the ACM International Conference on Parallel Architectures and Compilation Techniques, PACT ’20 (ACM, 2020).
- Harju et al. (2013) A. Harju, T. Siro, F. F. Canova, S. Hakala, and T. Rantalaiho, “Computational Physics on Graphics Processing Units,” in Lecture Notes in Computer Science (Springer Berlin Heidelberg, 2013) p. 3–26.
- Gao et al. (2006) X. Gao, S. Krishnamoorthy, S. K. Sahoo, C.-C. Lam, G. Baumgartner, J. Ramanujam, and P. Sadayappan, “Efficient Search-Space Pruning for Integrated Fusion and Tiling Transformations,” in Lecture Notes in Computer Science (Springer Berlin Heidelberg, 2006) p. 215–229.
- Sahoo et al. (2005) S. Sahoo, S. Krishnamoorthy, R. Panuganti, and P. Sadayappan, in ACM/IEEE SC 2005 Conference (SC’05) (IEEE, 2005).
- Yadav et al. (2024) R. Yadav, M. Bauer, D. Broman, M. Garland, A. Aiken, and F. Kjolstad, “Automatic Tracing in Task-Based Runtime Systems,” https://arxiv.org/abs/2406.18111 (2024), arXiv:2406.18111 [cs.DC] .
- Johnson et al. (2001) J. Johnson, R. W. Johnson, D. A. Padua, and J. Xiong, “Searching for the Best FFT Formulas with the SPL Compiler,” in Lecture Notes in Computer Science (Springer Berlin Heidelberg, 2001) p. 112–126.
- Meirom et al. (2022) E. Meirom, H. Maron, S. Mannor, and G. Chechik, in International Conference on Machine Learning (PMLR, 2022) pp. 15278–15292.
- Bondhugula et al. (2008) U. Bondhugula, A. Hartono, J. Ramanujam, and P. Sadayappan, ACM SIGPLAN Notices 43, 101–113 (2008).
- Lai et al. (2013) P.-W. Lai, K. Stock, S. Rajbhandari, S. Krishnamoorthy, and P. Sadayappan, in Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, SC13 (ACM, 2013).
- Psarras et al. (2022) C. Psarras, L. Karlsson, J. Li, and P. Bientinesi, “The landscape of software for tensor computations,” https://arxiv.org/abs/2103.13756 (2022), arXiv:2103.13756 [cs.MS] .
- Matthews (2018b) D. A. Matthews, SIAM Journal on Scientific Computing 40, C1 (2018b).
- Brabec et al. (2020) J. Brabec, J. Brandejs, K. Kowalski, S. Xantheas, O. Legeza, and L. Veis, Journal of Computational Chemistry 42, 534–544 (2020).
- Peng et al. (2016) C. Peng, J. A. Calvin, F. Pavošević, J. Zhang, and E. F. Valeev, The Journal of Physical Chemistry A 120, 10231 (2016).
- Levy, Solomonik, and Clark (2020) R. Levy, E. Solomonik, and B. K. Clark, in SC20: International Conference for High Performance Computing, Networking, Storage and Analysis (IEEE, 2020).
- Lyakh et al. (2022) D. I. Lyakh, T. Nguyen, D. Claudino, E. Dumitrescu, and A. J. McCaskey, Frontiers in Applied Mathematics and Statistics 8, 838601 (2022).
- Sharir, Shashua, and Carleo (2022) O. Sharir, A. Shashua, and G. Carleo, Physical Review B 106, 205136 (2022).
- Chi and Kolda (2012) E. C. Chi and T. G. Kolda, SIAM Journal on Matrix Analysis and Applications 33, 1272 (2012).
- Sidiropoulos et al. (2017) N. D. Sidiropoulos, L. D. Lathauwer, X. Fu, K. Huang, E. E. Papalexakis, and C. Faloutsos, IEEE Transactions on Signal Processing 65, 3551 (2017).
- Shi et al. (2016) Y. Shi, U. N. Niranjan, A. Anandkumar, and C. Cecka, in 2016 IEEE 23rd International Conference on High Performance Computing (HiPC) (IEEE, 2016).
- Georganas et al. (2021) E. Georganas, D. Kalamkar, S. Avancha, M. Adelman, C. Anderson, A. Breuer, J. Bruestle, N. Chaudhary, A. Kundu, D. Kutnick, F. Laub, V. Md, S. Misra, R. Mohanty, H. Pabst, B. Ziv, and A. Heinecke, in Proceedings of the International Conference for High Performance Computing, Networking, Storage and Analysis (ACM, 2021).
- Springer and Bientinesi (2017) P. Springer and P. Bientinesi, “Design of a high-performance GEMM-like Tensor-Tensor Multiplication,” https://arxiv.org/abs/1607.00145 (2017), arXiv:1607.00145 [cs.MS] .
- cuT (2019) “NVIDIA cuTENSOR: A High-Performance CUDA Library For Tensor Primitives,” https://docs.nvidia.com/cuda/cutensor/latest/index.html (2019).
- Bientinesi et al. (2022) P. Bientinesi, D. Ham, F. Huang, P. H. J. Kelly, P. S. Sadayappan, and E. Stow, Dagstuhl Reports , 1 (2022).
- Valeev (2023) E. Valeev, “Basic Programming of TiledArray,” https://valeevgroup.github.io/tiledarray/dox-master/_basic-_programming.html (2023), [Online; accessed 13-July-2023].
- cppreference.com community (2023) cppreference.com community, “C++ reference MDSPAN,” https://en.cppreference.com/w/cpp/container/mdspan (2023), [Online; accessed 13-July-2023].
- Murray et al. (2023) R. Murray, J. Demmel, M. W. Mahoney, N. B. Erichson, M. Melnichenko, O. A. Malik, L. Grigori, P. Luszczek, M. Dereziński, M. E. Lopes, T. Liang, H. Luo, and J. Dongarra, “Randomized Numerical Linear Algebra : A Perspective on the Field With an Eye to Software,” https://arxiv.org/abs/2302.11474 (2023), arXiv:2302.11474 [math.NA] .
- TAPP working group (2024) TAPP working group, “Tensor Algebra Processing Primitives,” https://github.com/TAPPorg/tensor-interfaces (2024).
- Pototschnig et al. (2021) J. V. Pototschnig, A. Papadopoulos, D. I. Lyakh, M. Repisky, L. Halbert, A. S. P. Gomes, H. J. Aa. Jensen, and L. Visscher, Journal of Chemical Theory and Computation 17, 5509 (2021).
- tal (2018) “TAL-SH,” https://github.com/DmitryLyakh/TAL_SH (2018).
- Mutlu et al. (2023) E. Mutlu, A. Panyala, N. Gawande, A. Bagusetty, J. Glabe, J. Kim, K. Kowalski, N. P. Bauman, B. Peng, H. Pathak, J. Brabec, and S. Krishnamoorthy, The Journal of Chemical Physics 159, 024801 (2023).
- Palmer et al. (2022) B. Palmer, E. Apra, A. Panyala, J. Hammond, C. Pe, and S. Krishnamoorthy, “GlobalArrays,” https://github.com/GlobalArrays/ga (2022).
- Zhang et al. (2020) T. Zhang, X.-Y. Liu, X. Wang, and A. Walid, IEEE Transactions on Parallel and Distributed Systems 31, 595 (2020).
- Millette, Ma, and Karunanidhi (2024) C. Millette, C. Ma, and M. Karunanidhi, “hipTensor,” https://github.com/ROCm/hipTensor (2024).
- Paszke et al. (2019) A. Paszke, S. Gross, F. Massa, A. Lerer, J. Bradbury, G. Chanan, T. Killeen, Z. Lin, N. Gimelshein, L. Antiga, A. Desmaison, A. Kopf, E. Yang, Z. DeVito, M. Raison, A. Tejani, S. Chilamkurthy, B. Steiner, L. Fang, J. Bai, and S. Chintala, in Advances in Neural Information Processing Systems 32 (Curran Associates, Inc., 2019) pp. 8024–8035.
- Abadi (2016) M. Abadi, in Proceedings of the 21st ACM SIGPLAN International Conference on Functional Programming (ACM, 2016).
- Crawford and Schaefer (2000) T. D. Crawford and H. F. Schaefer, Reviews in Computational Chemistry 14, 33 (2000).
- Kats (2024) D. Kats, “CCDiag - Coupled-Cluster diagrams in LaTeX,” https://github.com/fkfest/ccdiag (2024).
- Meurer et al. (2017) A. Meurer, C. P. Smith, M. Paprocki, O. Čertík, S. B. Kirpichev, M. Rocklin, A. Kumar, S. Ivanov, J. K. Moore, S. Singh, T. Rathnayake, S. Vig, B. E. Granger, R. P. Muller, F. Bonazzi, H. Gupta, S. Vats, F. Johansson, F. Pedregosa, M. J. Curry, A. R. Terrel, v. Roučka, A. Saboo, I. Fernando, S. Kulal, R. Cimrman, and A. Scopatz, PeerJ Computer Science 3, e103 (2017).
- Visscher et al. (2024) L. Visscher, H. J. Aa. Jensen, R. Bast, A. S. P. Gomes, T. Saue, I. A. Aucar, V. Bakken, J. Brandejs, C. Chibueze, J. Creutzberg, K. G. Dyall, S. Dubillard, U. Ekström, E. Eliav, T. Enevoldsen, E. Faßhauer, T. Fleig, O. Fossgaard, L. Halbert, E. D. Hedegård, T. Helgaker, B. Helmich-Paris, J. Henriksson, M. van Horn, M. Iliaš, C. R. Jacob, S. Knecht, S. Komorovský, O. Kullie, J. K. Lærdahl, C. V. Larsen, Y. S. Lee, N. H. List, H. S. Nataraj, M. K. Nayak, P. Norman, A. Nyvang, M. Olejniczak, J. Olsen, J. M. H. Olsen., A. Papadopoulos, Y. C. Park, J. K. Pedersen, M. Pernpointner, J. V. Pototschnig, R. Di Remigio Eikås, M. Repiský, K. Ruud, P. Salek, B. Schimmelpfennig, B. Senjean, A. Shee, J. Sikkema, A. Sunaga, J. Thyssen, J. van Stralen, M. L. Vidal, S. Villaume, O. Visser, T. Winther, S. Yamamoto, and X. Yuan, “DIRAC24,” https://zenodo.org/doi/10.5281/zenodo.10680560 (2024).
- Saue et al. (2020) T. Saue, R. Bast, A. S. P. Gomes, H. J. Aa. Jensen, L. Visscher, I. A. Aucar, R. Di Remigio, K. G. Dyall, E. Eliav, E. Fasshauer, T. Fleig, L. Halbert, E. D. Hedegård, B. Helmich-Paris, M. Iliaš, C. R. Jacob, S. Knecht, J. K. Laerdahl, M. L. Vidal, M. K. Nayak, M. Olejniczak, J. M. H. Olsen, M. Pernpointner, B. Senjean, A. Shee, A. Sunaga, and J. N. P. van Stralen, The Journal of Chemical Physics 152, 204104 (2020).
- NEC (2024) NEC, “NEC receives order for next-generation supercomputer system from Japan’s National Institutes for Quantum Science and Technology and National Institute for Fusion Science,” https://www.nec.com/en/press/202411/global_20241113_02.html (2024).
- Hewlett Packard Enterprise (2023) (HPE) Hewlett Packard Enterprise (HPE), “Cray MPICH,” https://www.mpich.org/static/docs/slides/2023-sc-bof/HPE.pdf (2023).
- Sucher (1980) J. Sucher, Phys. Rev. A 22, 348 (1980).
- Almoukhalalati et al. (2016) A. Almoukhalalati, S. Knecht, H. J. A. Jensen, K. G. Dyall, and T. Saue, The Journal of Chemical Physics 145, 074104 (2016).
- Jensen (2005) H. J. Aa. Jensen, (2005), Douglas–Kroll the Easy Way, Talk at Conference on Relativistic Effects in Heavy Elements - REHE, Mülheim, Germany, April, 2005. Available at https://doi.org/10.6084/m9.figshare.12046158.
- Kutzelnigg and Liu (2005) W. Kutzelnigg and W. Liu, J. Chem. Phys. 123, 241102 (2005).
- Iliaš and Saue (2007) M. Iliaš and T. Saue, J. Chem. Phys. 126, 064102 (2007).
- Liu and Peng (2009) W. Liu and D. Peng, J. Chem. Phys. 131, 031104 (2009).
- Hess et al. (1996) B. A. Hess, C. M. Marian, U. Wahlgren, and O. Gropen, Chem. Phys. Lett. 251, 365 (1996).
- (116) B. Schimmelpfennig, “AMFI, an atomic mean-field spin-orbit integral program,” University of Stockholm, Stockholm, Sweden; 1996.
- Knecht et al. (2022) S. Knecht, M. Repisky, H. J. Aa. Jensen, and T. Saue, The Journal of Chemical Physics 157, 114106 (2022).
- Dyall (2023) K. G. Dyall, “Dyall dz, tz, and qz basis sets for relativistic electronic structure calculations,” https://doi.org/10.5281/zenodo.7574629 (2023).
- Dunning (1989) T. H. Dunning, J. Chem. Phys. 90, 1007 (1989).
- Yuan, Visscher, and Gomes (2022) X. Yuan, L. Visscher, and A. S. P. Gomes, The Journal of Chemical Physics 156, 224108 (2022).
- Binkley, Pople, and Hehre (1980) J. S. Binkley, J. A. Pople, and W. J. Hehre, J. Am. Chem. Soc. 102, 939 (1980).
- Dill and Pople (1975) J. D. Dill and J. A. Pople, J. Chem. Phys. 62, 2921 (1975).
- Ditchfield, Hehre, and Pople (1971) R. Ditchfield, W. J. Hehre, and J. A. Pople, J. Chem. Phys. 54, 724 (1971).
- Kállay et al. (2020) M. Kállay, P. R. Nagy, D. Mester, Z. Rolik, G. Samu, J. Csontos, J. Csóka, P. B. Szabó, L. Gyevi-Nagy, B. Hégely, I. Ladjánszki, L. Szegedy, B. Ladóczki, K. Petrov, M. Farkas, P. D. Mezei, and Á. Ganyecz, J. Chem. Phys. 152, 074107 (2020).
- mrc (2024) “MRCC,” https://www.mrcc.hu/ (2024), MRCC, a quantum chemical program suite written by M. Kállay, P. R. Nagy, D. Mester, L. Gyevi-Nagy, J. Csóka, P. B. Szabó, Z. Rolik, G. Samu, J. Csontos, B. Hégely, Á. Ganyecz, I. Ladjánszki, L. Szegedy, B. Ladóczki, K. Petrov, M. Farkas, P. D. Mezei, and R. A. Horváth. See www.mrcc.hu.
- Kállay, Gauss, and Szalay (2003) M. Kállay, J. Gauss, and P. G. Szalay, The Journal of Chemical Physics 119, 2991–3004 (2003).
- Visscher, Lee, and Dyall (1996) L. Visscher, T. J. Lee, and K. G. Dyall, The Journal of Chemical Physics 105, 8769–8776 (1996).
- Gomes and Lyakh (2023) A. Gomes and D. I. Lyakh, “Relmbdev/mb-autogen: mb-autogen,” https://zenodo.org/record/8094430 (2023).
- Amdahl (2013) G. M. Amdahl, Computer 46, 38–46 (2013).
- Usov et al. (2024) D. P. Usov, Y. S. Kozhedub, V. V. Meshkov, A. V. Stolyarov, N. K. Dulaev, A. M. Ryzhkov, I. M. Savelyev, V. M. Shabaev, and I. I. Tupitsyn, Physical Review A 109, 042822 (2024).
- Koput (2024) J. Koput, Journal of Chemical Theory and Computation 20, 9041–9047 (2024).
- Raghavachari et al. (2013) K. Raghavachari, G. W. Trucks, J. A. Pople, and M. Head-Gordon, Chemical Physics Letters 589, 37–40 (2013).
- Kowalski and Piecuch (2000) K. Kowalski and P. Piecuch, The Journal of Chemical Physics 113, 18–35 (2000).
- Repisky et al. (2020) M. Repisky, S. Komorovsky, M. Kadek, L. Konecny, U. Ekström, E. Malkin, M. Kaupp, K. Ruud, O. L. Malkina, and V. G. Malkin, The Journal of Chemical Physics 152, 184101 (2020).